# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

### v1

@@ 4.1.1 Correctness and Complexity
PerpExplore-F2F solves the PerpExplore-BBH problem with $\lceil\log(n-1)\rceil+3$ agents having face to face communication and starting from the same node on a ring $R$ with $n$ nodes.
Proof 4.14.
We prove this theorem, by first ensuring that in any iteration of our Algorithm 1, every node is visited by at least one agent. Second, we show that, in any iteration there exists at least one alive agent.
Let $|S|$ be the length of the suspicious region and $k$ be the number of alive agents at the beginning of iteration $i$. Let $|S|=2$ and $k=2$, in this case we have, $S=\{S_{lft},S_{rgt}\}$. So, the two agents $a_{1}$ and $a_{2}$ (say) follows the sequence $SEQ0$, in which $a_{1}$ is instructed to visit $S_{lft}$ along clockwise direction, whereas $a_{2}$ is instructed to visit $S_{rgt}$ along counter-clockwise direction, while performing an iteration. This implies that, if $S_{lft}$ is at distance $l$ ($l>0$) in a clockwise direction, $a_{1}$ after traversing a distance of $l$ reaches $S_{lft}$, whereas $a_{2}$ traverses a distance of $n-l-1$ in counter-clockwise direction to visit $S_{rgt}$, in an iteration. This in turn shows that, the agents starting from $home$, visits each node of $R$ while performing an iteration. Note that, $|S|>2$ and $k=2$ never occurs by Corollary 4.10. So, the only possibility remains is that $|S|>2$ and $k>2$, in this situation, one among them say, $a_{1}$ executes the sequence $SEQ1$, another say, $a_{2}$ executes the sequence $SEQ2$, whereas the remaining agents does not leave $home$ while executing $SEQ3$, in an iteration. Observe that, $a_{1}$ traverses in a clockwise direction for $Dest$ distance, and reaches the destination node. Similarly, $a_{2}$ also reaches this same node, while traversing along a counter-clockwise direction for $n-Dest$ distance. This shows that, in any iteration, there exists two such agents, which explore the whole ring, hence exploration is guaranteed in each such iteration.
Next, we show that in every iteration, number of alive agents is at least 1. Let us consider during the $i$-th iteration, all the agents gets consumed by the byzantine black hole.
Note that, at the beginning of the $i-$th iteration, number of alive agent can not be 1. By contradiction, let us suppose there is only one agent alive at the beginning of $i-$th iteration which gets consumed by the byzantine black hole during this iteration. This implies $|S|\geq 2$ at the beginning of this iteration. Now by Lemma 4.11 and Corollary 4.10, there must exists at least two alive agents at the beginning of the $i-$th iteration which contradicts our assumption of one alive agent. Also observe that number of alive agent at the start of iteration $i$, can not be strictly greater than 2. Otherwise, let number of alive agents $k>2$ at the beginning of iteration $i$. Since during iteration $i$ all the $k$ agents are destroyed, this implies $|S|\geq 2$ Now since $|S|\geq 2$ and $k>2$, our algorithm ensures that, in the worst case during this iteration at most two agents can be consumed by the byzantine black hole. So after the iteration number of alive agents is at least $k-2\geq 1$. This contradicts our assumption that all agents are destroyed at the end of this iteration. Now the only remaining case is that exactly two agents are alive at the beginning of $i-$th iteration. Similarly as both of the agents are destroyed during the $i-$th iteration $|S|\geq 2$ at the beginning of this iteration. Note that $|S|$ will be exactly 2, for $k=2$, as by Corollary 4.10, we have if $|S|>2$ then $k$ must be at least 3. Now according to the algorithm, both the agents follows $SEQ0$, which instructs them to follow disjoint paths. Hence, this concludes that during this iteration, only can be consumed by the black hole, while there exists another agent which is alive even after the iteration, contradicting our assumption. Hence for all iterations number of alive agents must be greater or equals to 1. Thus, algorithm PerpExplore-F2F solves the PerpExplore-BBH problem with $\lceil\log(n-1)\rceil+3$ agents having face to face communication and starting from the same node on a ring $R$ with $n$ nodes.
4.2 Communication: Pebble
In this section, we consider the model, where each agent has an access to a movable token (termed as pebble) which can be carried by an agent from one node to another. This pebble acts as a mode of communication for the agents, as the agents can perceive the presence of a pebble at the current node. Moreover, an agent can also perceive the presence of other agents which are co-located at the current round (i.e., gather the Ids of the other co-located agents). Next, we discuss the idea of our algorithm PerpExplore-Coloc-Pbl (Algorithm 2) which solves the PerpExploration-BBH problem with 3 agents, each accompanying a pebble.
Description of PerpExplore-Coloc-Pbl: Given a ring $R$ with a byzantine black hole node $v_{b}$, a set of agents $A=\{a_{1},a_{2},a_{3}\}$ (where w.l.o.g ID of $a_{1}\leavevmode\nobreak\ <$ ID of $a_{2}\leavevmode\nobreak\ <$ ID of $a_{3}$), are initially co-located at a safe node, which we term as $home$. In this case as well, the agents have no knowledge about the position of the node $v_{b}$, the only set of knowledge the agents have are: the total number of nodes in the underlying topology $R$ (i.e., $n$), and also the knowledge that the initial node, i.e., $home$ is safe. So, the remaining arc of $n-1$ nodes in $R$ is a suspicious region (which we again term as $S$, where the cardinality of this arc is defined by $|S|$) for each of the three agents.
[…]

### v2

@@ 4.1 Pebble Model of Communication
Algorithm PerpExplore-Coloc-Pbl solves PerpExploration-BBH on a ring $R$ with 3 co-located and synchronous agents under the pebble model of communication with the presence of two pebbles initially co-located with the agents.
We first prove that if no agents are destroyed then either the agents continue with the exploration of ring $R$ without knowing the exact location of the Byzantine black hole, or there exists at least one agent that knows the exact location of the Byzantine black hole which can then explore the ring indefinitely avoiding the Byzantine black hole (See Lemma 9.1 in Appendix). Next we proved that, if one agent is destroyed while exploring the ring, then either there exists one agent that identifies the Byzantine black hole uniquely and continues to indefinitely explore all nodes of the ring, except the Byzantine black hole or the exploration continues while the length of the suspicious region (i.e., $|S|$) decreases to 2 from $n-1$ (Lemma 9.3 in Appendix). For the later case we proved that the exploration continues until another agent is destroyed by the Byzantine black hole (Remark 9.5 in Appendix). In this case when two agents are destroyed by the Byzantine black hole, the only alive agent can uniquely identify the Byzantine black hole without moving on to it (Lemma 9.6 in Appendix). Thus it can then perform the exploration of $R$ avoiding the Byzantine black hole. This proves Theorem 4.2.
4.2 Face-to-Face Model of Communication
In this section, we consider the Face-to-Face (also termed as F2F) model and prove the following theorem.
Theorem 4.3.
A team of 5 co-located and synchronous agents are sufficient to solve PerpExploration-BBH on a ring $R$ with $n$ nodes under the F2F model of communication.
In order to prove Theorem 4.3, we discuss an algorithm with 5 co-located agents, where each agent can communicate among themselves at the same node at the same round (i.e., each agent have F2F model of communication). Initially 3 lowest ID agents, say, $a_{1}$, $a_{2}$ and $a_{3}$ are chosen, where fourth lowest ID agent say, $a_{4}$, is associated with $a_{1}$ and the largest ID agent, say $a_{5}$, is associated with $a_{2}$. The idea of the algorithm resembles to that of the algorithm PerpExplore-Coloc-Pbl. Note that as in this case there is no existence of pebbles, hence, we configure the behaviour of the agents $a_{4}$ and $a_{5}$ in such a way that they act as pebbles, associated with $a_{1}$ and $a_{2}$ in PerpExplore-Coloc-Pbl, respectively. In PerpExplore-Coloc-Pbl when we say that an agent $a_{i}$ carries a pebble $p$, in this case we mean that the agent $a_{i}$ communicates a message carry with its associated agent $a_{i^{\prime}}$ such that both these agents simultaneously move together until further new instruction is communicated. On the contrary as per PerpExplore-Coloc-Pbl when we say that an agent $a_{i}$ drops a pebble $p$ at a node $v$, then in this case we mean that $a_{i}$ communicates a message drop with $a_{i^{\prime}}$, which in turn instructs $a_{i^{\prime}}$ not to move further and remain stationary at the node $v$ until further instruction is communicated. Hence, with this terminology, a team of 5 agents can execute the algorithm PerpExplore-Coloc-Pbl, and the correctness also follows similarly. This shows that a team of 5 co-located and synchronous agents are sufficient to solve PerpExploration-BBH under F2F model of communication. This proves Theorem 4.3.
4.3 Whiteboard Model of Communication
The algorithm PerpExplore-Coloc-Pbl can be simulated in whiteboard model of communication. A pebble on a node can be simulated by a bit of information, which is marked on the whiteboard of that node. Note that, collecting a pebble from the node is simulated by erasing the same bit of information, on that node and dropping the pebble can be simulated by marking the node with a bit of information as well on the whiteboard of that node. From this we get the following result for whiteboard model of communication.
Theorem 4.4.
A team of 3 synchronous co-located agents are necessary and sufficient to solve PerpExploration-BBH on a ring $R$ if each node are equipped with a whiteboard having constant memory.
Previously in [2], for asynchronous scheduler the tight bound for number of agents to solve this problem was 4. Now from Theorem 4.4, we get a trade-off between reducing the optimal number of agents required vs the scheduler, in presence of a more generalized version of gray hole as well (i.e., Byzantine black hole).
5 Perpetual Exploration with Scattered Agents
Here in this section, we discuss the problem of PerpExploration-BBH, when the agents are initially scattered on more than one nodes, where each of these starting nodes are assumed to be safe. We investigate this problem, under different communication models. First, we discuss the pebble model of communication and subsequently we discuss the whiteboard model of communication.
5.1 Pebble Model of Communication
Our goal in this section is to prove the following theorem.
Theorem 5.1.
A team of 4 synchronous scattered agents with one pebble each is necessary and sufficient to solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes, when the agents are initially scattered on $R$.
The necessary part follows from Corollary 3.5. For the sufficient part, we present an algorithm PerpExplore-Scat-Pbl, that can solve the PerpExploration-BBH problem in the above context. In this algorithm we assumed that 4 agents are initially scattered at 4 different nodes of $R$. To include the remaining cases where the agents are initially scattered at less than or equal to 3 nodes, a slight modification of PerpExplore-Scat-Pbl is enough which we describe in Remark 10.7 in Appendix. Here we describe only the main idea of the algorithm. The detailed description and pseudocodes of the algorithm can also be found at the Appendix (Section 10.1).
Brief Idea of the Algorithm: Let us consider, the starting position of $a_{0}$ to be the first, after which the starting position of $a_{1},a_{2}$ and $a_{3}$, respectively follows in clockwise order. Let $h_{i}$ be the starting node of an agent $a_{i}$ (i.e., $home$ of $a_{i}$), where $i\in\{0,1,2,3\}$. By $Seg(a_{i})$ we define the clockwise arc starting from the node $h_{i}$ and ending at $h_{(i+1)\pmod{4}}$ (called segment of $a_{i}$). If no agents are destroyed by the Byzantine black hole $v_{b}$ then the exploration of $R$ goes as follows.
Agent $a_{i}$ moves clockwise along $Seg(a_{i})$ leaving its pebble at $h_{i}$ until it reaches $h_{(i+1)\pmod{4}}$ (i.e., the end of $Seg(a_{i})$). An agent can distinguish the node $h_{(i+1)\pmod{4}}$ by seeing the pebble left there by the agent $a_{(i+1)\pmod{4}}$. When $a_{i}$ reaches $h_{(i+1)\pmod{4}}$ traversing $Seg(a_{i})$ for the first time, it knows the length of the segment $Seg(a_{i})$ and stores the length in its own memory. For further traversals it does not depend on seeing pebbles at the end nodes of the segment. It can simply use the length of the segment. After $a_{i}$ reaches $h_{(i+1)\pmod{4}}$, it waits there for a certain number of rounds so that the other agents can in the meantime reach the endpoints of there corresponding segments, while moving along a clockwise direction. After this waiting, $a_{i}$ collects the pebble (if exists) from $h_{(i+1)\pmod{4}}$ (the pebble left by $a_{(i+1)\pmod{4}}$) and starts moving counter-clockwise along with the collected pebble until it reaches its own $home$, i.e., $h_{i}$ along with the pebble it was carrying. The waiting time at $h_{(i+1)\pmod{4}}$ also ensures that all agents starts moving counter-clockwise at the same round. Also note that, if $a_{i}$ does not find any pebble to collect at $h_{(i+1)\pmod{4}}$ it starts moving counter-clockwise without any pebble (this case can only happen if $a_{(i+1)\pmod{4}}$ is destroyed at the Byzantine black hole node $v_{b}$, while it was returning back after collecting the pebble from $h_{(i+2)\pmod{4}}$ in some previous round). After $a_{i}$ reaches $h_{i}$ moving counter-clockwise, it again waits for another set of rounds before it repeats the whole process again. Again, the waiting time at $h_{i}$ is configured in such a way that, all agents start repeating the process at the same round (for details on the precise value of these waiting times, refer to the detailed description of the algorithm in Appendix in Section 10.1). Note that, since $\cup_{i=0}^{3}Seg(a_{i})=R$, if no agents are destroyed, the perpetual exploration of $R$ happens.
Figure 2: (a) $h_{i}$’s are $home$ marked as green. Agent $a_{i}$ is on $h_{i}$ initially with a pebble. We name the pebble initially at $h_{i}$ as $p_{i}$, but in reality they are anonymous. (a-b) Each agent moves clockwise until the next $home$ (i.e., $h_{(i+1)\pmod{4}}$) without carrying any pebble. The agents already reached ( here $a_{1}$ and $a_{3}$) waits for others. (c-e) All agents are on their clockwise nearest $home$. They start moving counter clockwise together with the pebble present at their current location towards their initial $home$. The agents which already reaches their $home$, wait for others to reach their $home$.
The exploration can be hampered only if, an agent is destroyed at the Byzantine black hole node $v_{b}$. Note that, since $Seg(a_{i})\cap Seg(a_{j})$ is either empty or consists of a safe node and in addition to that, as a segment is explored by only one agent. Hence, at most one agent can be destroyed while exploring their respective segments, as described above. Without loss of generality, let $v_{b}\in Seg(a_{j})$, for some $j\in\{0,1,2,3\}$. Now there are two cases.
Case-I: Let $a_{j}$ is destroyed while moving clockwise. For this case, $a_{j}$ fails to collect the pebble at $h_{(j+1)\pmod{4}}$ left by $a_{(j+1)\pmod{4}}$ and return $h_{j}$. So, when $a_{(j+1)\pmod{4}}$, returns $h_{(j+1)\pmod{4}}$ after moving counter-clockwise along with the pebble it has collected from $h_{(j+2)\pmod{4}}$, it finds two pebble at $h_{(j+1)\pmod{4}}$. This is considered by $a_{(j+1)\pmod{4}}$ as an anomaly and it learns that $v_{b}$ must be in the arc, which is in the counter-clockwise direction starting from $h_{(j+1)\pmod{4}}$. In this scenario, $a_{(j+1)\pmod{4}}$ waits a certain number of rounds (if needed) to ensure that every other alive agents have reached their corresponding $home$. Then $a_{(j+1)\pmod{4}}$ starts moving clockwise with both the pebbles. The aim of this move is to meet and gather with the other alive agents. Note that the other alive agents wait at their corresponding $home$, after moving along the counter-clockwise direction. It is because, at their respective $home$ they do not detect any anomaly. The waiting time is provided in such a way that it is enough for $a_{(j+1)\pmod{4}}$ to detect anomaly and meet both the remaining alive agents after moving clockwise, while the other agents are waiting at their $home$. The agent $a_{(j+1)\pmod{4}}$ first meets $a_{(j+2)\pmod{4}}$ at $h_{(j+2)\pmod{4}}$ while it moves clockwise after detecting anomaly. Then both of them moves together, while $a_{(j+2)\pmod{4}}$ carries the pebble, which it was earlier carrying back to $home$. They move until they meet with $a_{(j+3)\pmod{4}}$ at $h_{(j+3)\pmod{4}}$. Note that at this moment 3 agents are at a node (which is $h_{(j+3)\pmod{4}}$), with at least 3 pebbles (one carried by $a_{(j+3)\pmod{4}}$ and two carried by $a_{(j+2)\pmod{4}}$). In this case they execute the algorithm PerpExplore-Coloc-Pbl and achieve PerpExploration-BBH of the ring $R$.
Case-II: Let $a_{j}$ be destroyed at $v_{b}$ while it was moving counter-clockwise along with the pebble it collected from $h_{(j+1)\pmod{4}}$. Then in this case, when all alive agents return to their corresponding $home$, none of them finds any anomaly as all agent sees exactly one pebble at their $home$. So, they wait and start the exploration again at the same round. Note that in this scenario, all agents except $a_{j}$, move clockwise to the end points of their corresponding segments, waits there and collects pebble (if any) and moves back to their corresponding $home$ again. Now, since $a_{j}$ was destroyed earlier, the pebble left by $a_{(j+1)\pmod{4}}$ was picked by no agents and thus, while $a_{(j+1)\pmod{4}}$ returns back to $h_{(j+1)\pmod{4}}$ it finds two pebbles and does the same execution as explained in Case-I.
The basic idea of this algorithm is to gather three agents with at least 3 pebbles (refer Remark 10.1) in the expense of one agent and then execute PerpExplore-Coloc-Pbl to solve the PerpExplroation-BBH problem.
Sketch of Correctness: To proof the correctness of the algorithm PerpExplore-Scat-Pbl, we have to prove the following theorem. Here we only provide the proof idea. For details see Appendix (Section 10.2).
Theorem 5.2.
Algorithm PerpExplore-Scat-Pbl solves PerpExploration-BBH problem of a ring $R$ with 4 synchronous and scattered agents under the pebble model of communication where each agents are equipped with a pebble.
First in Corollary 10.4, we proved that algorithm PerpExplore-Scat-Pbl guarantees exploration if no agents are destroyed. This corollary is a direct consequence of the Lemma 10.2. Further we showed that if an agent is destroyed then exactly one agent detects anomaly and that agent gathers with the remaining two alive agents within finite rounds (Lemma 10.5 in Appendix). The rest of the proof follows from Theorem 4.2.
5.2 Whiteboard Model of Communication
The main aim of this section is to prove the following theorem.
Theorem 5.3.
A team of 3 synchronous agents are necessary and sufficient to solve the problem PerpExploration-BBH on a ring $R$ with $n$ nodes, when each node of $R$ has a whiteboard of $O(\log n)$ bits of memory, irrespective of their starting location.
The necessary part follows from Corollary 3.2. The sufficiency part for co-located initial situation follows from Theorem 4.4. For the other cases where the agents are initially scattered at more than one starting nodes, we propose an algorithm PerpExplore-Scat-Whitbrd. This algorithm is designed for the case when all the agents are starting at different nodes. Note that, this algorithm can be easily modified a bit to include the case where initially three agents are scattered at two distinct nodes (refer the modification in Appendix, Section 11.3). For detailed description and the pseudo code, see Appendix (Section 11.1).
Brief Description of the Algorithm: Let $a_{0},a_{1}$ and $a_{2}$ be three agents starting from the nodes $h_{0},h_{1}$ and $h_{2}$, respectively, where these nodes are in a clockwise order. Let $Seg(a_{i})$ (also called ‘Segment of $a_{i}$’) be the clockwise arc starting from $h_{i}$ and ending at $h_{(i+1)\pmod{3}}$. The algorithm first ensures that the ring is explored perpetually if no agents are destroyed. In order to do this, the algorithm goes as follows.
An agent, say $a_{i}$, first erases all previously stored data (if any) at $h_{i}$. Then it writes the message (home, ID($a_{i}$)) at $h_{i}$ and starts moving clockwise. This type of message is called a home type message that indicates it is $home$ of $a_{i}$. $a_{i}$ moves clockwise until it reaches $h_{(i+1)\pmod{3}}$. It distinguishes $h_{(i+1)\pmod{3}}$ by seeing the home type message left by $a_{(i+1)\pmod{3}}$. When $a_{i}$ moves clockwise, it also marks each node of $Seg(a_{i})$, except $h_{i}$ and $h_{(i+1)\pmod{3}}$, by writing right, after erasing any previous such markings (if at all exists) at each such nodes. After $a_{i}$ reaches $h_{(i+1)\pmod{3}}$ it waits for a certain number of rounds (if needed), so that the other agents (say $a_{j}$) gets enough time to reach endpoints of their corresponding segment (i.e.,$h_{(j+1)\pmod{3}}$). After this waiting, each alive agent $a_{i}$ write the message (visited, ID($a_{i}$)) at $h_{(i+1)\pmod{3}}$ at the same round. This type of messages is termed as a visited type message, which indicates that an agent with its corresponding ID, also mentioned in the visited type message, has visited the respective node. Then each $a_{i}$ waits for $n$ rounds at $h_{(i+1)\pmod{3}}$ and then starts moving counter-clockwise at the same round. $a_{i}$ moves counter-clockwise until it reaches $h_{i}$, i.e., its own $home$. While moving counter-clockwise, $a_{i}$ erases previously written right marking (which it marked while moving along clockwise direction) from each nodes of $Seg(a_{i})$ and writes left there upon arriving (except at $h_{i}$ and at $h_{(i+1)\pmod{3}}$). When $a_{i}$ reaches its own $home$ (i.e., $h_{i}$), it waits again for a certain number of rounds, upon seeing the visited type message, left there by $a_{(i-1)\pmod{3}}$. This waiting period is enough for each of the other alive agents to reach their corresponding $home$. After this waiting period is over, all of the agents, starts repeating the same procedure again together, from the same round. This procedure ensures that, if no agent is destroyed at the Byzantine black hole node $v_{b}$, then the perpetual exploration of $R$ continues. This can only be hampered, only if an agent gets destroyed at the node $v_{b}$. Without loss of generality let $v_{b}\in Seg(a_{j})$ for some, $j\in\{0,1,2\}$. So only $a_{j}$ can be destroyed at $v_{b}$, while performing this exploration. It is because $a_{j}$ is the only agent to visit each node $u$ of $Seg(a_{j})$, where $u\in Seg(a_{j})\backslash\{h_{j},h_{(j+1)\pmod{3}}\}$. There can be two cases.
Case-I: Let $a_{j}$ be destroyed while it is moving in clockwise direction along $Seg(a_{j})$. This implies it fails to reach $h_{(j+1)\pmod{3}}$ and also fails to write a visited type message there. So, when $a_{(j+1)\pmod{3}}$ returns to its $home$ (i.e., $h_{(j+1)\pmod{3}}$), it finds no visited type message (as it should’ve, if $a_{i}$ was not destroyed). $a_{(j+1)\pmod{3}}$ interprets from this that, the segment which is adjacent to its own segment in counter-clockwise direction has the Byzantine black hole and an agent must have been destroyed there while moving clockwise. This information triggers it to move clockwise with the aim of gathering with the other agent (i.e., $a_{(j+2)\pmod{3}}$). Note that, when $a_{(j+1)\pmod{3}}$ starts moving, the other agent is waiting and the waiting time is sufficient for the moving agent to meet it. After they meet, $a_{(j+1)\pmod{3}}$ shares the information about its direction of movement in the whiteboard of $h_{(j+2)\pmod{3}}$. With this, they again start moving clockwise together until they reach $h_{j}$. The agents can distinguish $h_{j}$ by the home type message written there by $a_{j}$ before it was destroyed. Note that in the clockwise direction each node after $h_{j}$ up to the previous node of $v_{b}$ were marked right by $a_{j}$ before it was destroyed. So from $h_{j}$, $a_{(j+1)\pmod{3}}$ and $a_{(j+2)\pmod{3}}$ starts moving cautiously clockwise. That is, while both agents ($a_{L}$ and $a_{H}$, where $a_{L}$ and $a_{H}$ are the agent with lowest ID and highest ID among $a_{(j+1)\pmod{3}}$ and $a_{(j+2)\pmod{3}}$ respectively) are at a node $v_{0}$, $a_{L}$ moves clockwise to the next node, say $v_{1}$ while the other agent waits at $v_{0}$. If at $v_{1}$, $a_{L}$ sees the right marking, it interprets $v_{1}$ is safe. So, it comes back to $v_{0}$. At $v_{0}$, seeing that $a_{L}$ has returned, $a_{H}$ also interprets $v_{1}$ to be safe. So, in the next round both $a_{L}$ and $a_{H}$ moves to $v_{1}$ together and repeats the process from $v_{1}$ again. When $a_{L}$ reaches $v_{b}$, it sees no right marking there. $a_{L}$ if alive, interprets this by determining the current node to be Byzantine black hole. So, it starts perpetual exploration of $R$, except $v_{b}$. Otherwise, if $a_{L}$ gets destroyed at $v_{b}$, it does not return to the previous node where $a_{H}$ is waiting. Even after waiting, when $a_{H}$ sees $a_{L}$ has not returned, it interprets this incident as the next clockwise node is the Byzantine black hole and starts exploring the ring $R$ avoiding that node. Thus for this case the perpetual exploration continues.
Case-II: Let $a_{j}$ be destroyed at $v_{b}$ while it is moving counter-clockwise. In this case both the alive agents $a_{(j+1)\pmod{3}}$ and $a_{(j+2)\pmod{3}}$ find visited type message after returning to their corresponding $home$. This is because, $a_{j}$ is destroyed at $v_{b}$ after writing visited type message at $h_{(j+1)\pmod{3}}$. So, all alive agents after waiting, starts repeating the exploration procedure again. The agents first erase the whiteboard content at their corresponding $home$ and then writes the corresponding home type message before it starts moving clockwise until the end of their corresponding segment. Note that since $a_{j}$ was destroyed earlier (even before the repetition starts), $a_{(j-1)\pmod{3}}$ finds the visited type message which was written there earlier by it, is still present (which was supposed to be erased if $a_{j}$ was alive). From this information, $a_{(j-1)\pmod{3}}$ interprets that, $v_{b}$ must be in the segment adjacent to its own segment in the clockwise direction. It also interprets that, an agent must have been destroyed at $v_{b}$ while it was moving counter-clockwise. This is because, if the agent was destroyed while it was moving clockwise then it would have been detected by the other agents when they are at their $home$ (as described in Case-I), which is before the start of the next clockwise move (note that the agents have already started next clockwise move). So, this information triggers $a_{(j-1)\pmod{3}}$ to move counter-clockwise with the aim to gather with $a_{(j-2)\pmod{3}}$. Since $a_{(j-1)\pmod{3}}$ starts its move while $a_{(j-2)\pmod{3}}$ waits at $home$ and also, since the waiting time is sufficient, $a_{(j-1)\pmod{3}}$ meets $a_{(j-2)\pmod{3}}$ during the waiting period at $h_{(j-1)\pmod{3}}$. So, after they meet, $a_{(j-1)\pmod{3}}$, communicates the direction of its move to $a_{(j-2)\pmod{3}}$ on the whiteboard of $h_{(j-1)\pmod{3}}$ and they move together until they reach $h_{(j-2)\pmod{3}}(=h_{(j+1)\pmod{3}})$ together. They distinguish the node by the home type message left there by $a_{(j-2)\pmod{3}}$ before the start of moving clockwise. Note that, in the counter-clockwise direction, the nodes in $Seg(a_{j})$, starting from the next node of $h_{(j+1)\pmod{3}}$ up to the previous node of $v_{b}$ are the only nodes those were marked left by $a_{j}$ before it was destroyed. So, from $h_{(j+1)\pmod{3}}$, the alive agents $a_{(j+1)\pmod{3}}$ and $a_{(j+2)\pmod{3}}$ start moving cautiously in the counter-clockwise direction. Among $a_{(j+1)\pmod{3}}$ and $a_{(j+2)\pmod{3}}$, an agent with lowest ID is denoted by $a_{L}$ and the agent with highest ID is denoted by $a_{H}$. The cautious walk is same as described in Case-I, except for the fact that here after moving one node in the counter-clockwise direction, the agent $a_{L}$ searches for the left marking. If it finds such marking it returns to $a_{H}$ and then both moves counter-clockwise and repeats the process. On the other hand if $a_{L}$ does not find any left marking (it must be $v_{b}$) and it stays alive, $a_{L}$ moves out of that node and starts exploring $R$ avoiding that node. On the contrary, if $a_{L}$ gets destroyed then $a_{H}$ sees that $a_{L}$ has not returned while it should have. From this incident it interprets next counter-clockwise node is $v_{b}$ and it starts exploring $R$ avoiding that node.
Sketch of Correctness: Here we give a brief glimpse on how we proved the following theorem. For details see Appendix (Section 11.2)
Theorem 5.4.
Algorithm PerpExplore-Scat-Whitbrd solves PerpExploration-BBH problem of a ring $R$ with $n$ nodes and with 3 synchronous agents initially scattered under the whiteboard model of communication
In order to prove this theorem, we first define a cautious start node. Let $a_{i}$ be the agent destroyed at the Byzantine black hole first. We define cautious start node to be $h_{i}$, if $a_{i}$ was moving clockwise when destroyed, otherwise it is $h_{(i+1)\pmod{3}}$. We first prove that after one agent is destroyed, the remaining two agents gather at the cautious start node. This result can be found in details in Lemma 11.11 in Appendix. Note that $a_{i}$ marks each node on the arc between the cautious start node and the Byzantine black hole with either left or right markings depending on the direction it was moving before it was destroyed. We proved that the alive agent can know the direction of $a_{i}$ before it was destroyed. This can be found in details in Lemma 11.1 and Lemma 11.3 in Appendix. After the alive agents reach the cautious start node, they start moving cautiously looking for the marking based on the direction of $a_{i}$ before it was destroyed. We proved that at least an agent among the two moving cautiously will stay alive knowing the exact location of the Byzantine black hole (Lemma 11.13 in Appendix). Hence this agent can explore $R$ perpetually avoiding the Byzantine black hole. This proves Theorem 5.4.
6 Conclusion
The paper addresses perpetual exploration of a ring network in the presence of a malicious node, which we call as a Byzantine black hole. This problem, termed as PerpExploration-BBH, is explored under three communication models (Face-To-Face, Pebble, and Whiteboard) considering various initial scenarios (co-located or scattered agents) with the aim of minimizing number of agents. We proposed optimal results (in terms of number of agents), for Pebble and Whiteboard communication models, under both initial scenarios. Further, an upper bound of 5 agents and a lower bound of 3 agents is provided for Face-to-Face model, in case of co-located agents.
Future research could focus on proposing an optimal bound for Face-To-Face co-located scenario, whereas proposing constructive lower and upper bounds for Face-To-Face scattered scenario. Additionally, investigating this problem in different scheduler models can be another way of future direction.
Appendix
7 Proof of Theorem 3.1
Statement: A set of two synchronous agents in a ring $R$ of size $n$ cannot solve
PerpExploration-BBH, even in presence of whiteboard if number of possible
consecutive Byzantine black hole positions is at least 3.
Proof: Let $v_{1},v_{2}$ and $v_{3}$ be three possible consecutive Byzantine black hole positions in a ring $R$. Let two agents $a_{1}$ and $a_{2}$ be sufficient to solve the PerpExploration-BBH problem on $R$. Thus there exists an algorithm $\mathcal{A}$ such that $a_{1}$ and $a_{2}$ can solve PerpExploration-BBH by executing $\mathcal{A}$. Without loss of generality let $a_{1}$ explores $v_{2}$ and it moves to $v_{2}$ for the first time at round $t$ from vertex $v_{1}$. Then $a_{1}$ must have been at the vertex $v_{1}$ at round $t-1$. Let us take two copies $R_{1}$ and $R_{2}$ of the same ring $R$. In $R_{1}$, $v_{1}$ is the Byzantine black hole and in $R_{2}$, $v_{2}$ is the Byzantine black hole. Let the adversary in $R_{1}$ destroy $a_{1}$ at round $t-1$ and in $R_{2}$ destroy $a_{1}$ at round $t$. We claim that at round $t$ and $t-1$, $a_{2}$ can not be on either of $v_{1}$ or $v_{2}$. Note that at $t$-th round $a_{2}$ can not be on $v_{2}$ otherwise, both $a_{1}$ and $a_{2}$ get destroyed at the same round and no other agent is left to explore further. Similarly, at round $t-1$, $a_{2}$ can not be on $v_{1}$. Now consider the case when at $t-$ th round $a_{2}$ is at $v_{1}$. Since at round $t-1$, $a_{2}$ can not be on $v_{2}$ there can be two cases either at round $t-1$, $a_{2}$ was in $u_{0}$ ($u_{0}$ is another adjacent node of $v_{1}$ except $v_{2}$) or was in $v_{1}$ itself. We first argue that at round $t-1$, $a_{2}$ can not be at $u_{0}$. Otherwise, since the whiteboard content at nodes except at $v_{1}$ and $v_{2}$ are the same in $R_{1}$ and $R_{2}$ at round $t$ (due to the same execution by $a_{1}$ and $a_{2}$ up to round $t-1$ and at round $t$ for $a_{2}$ in both $R_{1}$ and $R_{2}$). Now since at round $t$, all previous data on the whiteboard except $v_{1}$ and $v_{2}$ does not help $a_{2}$ to distinguish between $R_{1}$ and $R_{2}$, the problem now can be thought of solving PerpExploration-BBH with one agent where the number of possible consecutive Byzantine black hole position is atleast two. Now this is impossible to solve. So, at $t-1$-th round $a_{2}$ can not be on $u_{0}$. We now only have to prove that $a_{2}$ can not be at $v_{2}$ during round $t-1$. Let in $R_{2}$, adversary activates the Byzantine black hole at $v_{2}$ at round $t-1$ which destroys $a_{2}$ at round $t-1$. Now since the effect of this destruction of agent $a_{2}$ does not affect the inputs of $a_{1}$ at round $t-1$ on $v_{1}$, it moves to $v_{2}$ at round $t$ where the Byzantine black hole is activated again by adversary destroying $a_{1}$ too. Thus $a_{2}$ must remain outside of $v_{1}$ and $v_{2}$ during round $t$ and $t-1$ while executing $\mathcal{A}$. So, at round $t$ since all whiteboard content are same for $R_{1}$ and $R_{2}$ (except for nodes $v_{1}$ and $v_{2}$ whose content only differs after round $t-2$ for $R_{1}$ and $R_{2}$ which can not be known by $a_{2}$ as it was not there during that rounds), they can not help $a_{2}$ to distinguish between $R_{1}$ and $R_{2}$. So the problem now can be thought of as solving PerpExploration-BBH with one agent where the number of possible consecutive Byzantine black hole positions is at least 2. And this is impossible. So there doesn’t exist any algorithm that solves PerpExploration-BBH with two agents on a Ring $R$ where the number of possible consecutive positions of the Byzantine black hole is at least 3.
8 Proof of Theorem 3.4
[…]